OpenStudy (anonymous):

Factor completely: 3x^2 + 5x + 1 a. (3x + 1)(x + 1) b. (3x + 5)(x + 1) c. (3x − 5)(x + 1) d. Prime

Parth (parthkohli):

Didn't I just teach you?

OpenStudy (ajprincess):

Prime. It cannot be factorised.

Parth (parthkohli):

Correct @ajprincess

OpenStudy (anonymous):

I'm working on it. 3*1 = 3. 3+2 = 5.

Parth (parthkohli):

But you can't do that 3 * 2 = 6!

OpenStudy (ajprincess):

Thanx @ParthKohli

OpenStudy (anonymous):

Ohhh, I'm glad I stopped where I was then.

OpenStudy (anonymous):

so what if b has an x?

Parth (parthkohli):

If you can't factor it then it cant be factored ;P

OpenStudy (anonymous):

like this, for example. 3x2 − x − 4.

OpenStudy (anonymous):

3x^2 - x - 4.

Parth (parthkohli):

You can factor this.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

b = -1?

OpenStudy (anonymous):

-12*

Parth (parthkohli):

-4 * 3 = -12 -4 * 3 = -12 It can be done

Parth (parthkohli):

Don't confuse it with a and b and c now. That was just a demonstration.

OpenStudy (anonymous):

What does b = though?

OpenStudy (ajprincess):

-1

Parth (parthkohli):

Just see the first and last terms and see their coefficient, multiply them. Then split the middle term such that the split coefficients have the same product.

OpenStudy (anonymous):

a = 3. b = ? c = -4.

Parth (parthkohli):

b = -1

Parth (parthkohli):

-x = -1x so that's very very clear

OpenStudy (anonymous):

-4 * 3 = 12. So, I have -1x and 12.

Parth (parthkohli):

And -4 * 3 = 12 -4 + 3 = -1 :D

OpenStudy (anonymous):

(3x − 4)(x + 1)?

OpenStudy (anonymous):

I just don't get it.

Parth (parthkohli):

Split like this. \(\Large \color{Black}{\Rightarrow 3x^2 - 4x + 3x - 4 }\) \(\Large \color{Black}{\Rightarrow x(3x - 4) + 1(3x - 4) }\) \(\Large \color{Black}{\Rightarrow (x + 1)(3x - 4) }\) Correct :D

OpenStudy (ajprincess):

Good work @ParthKohli.:)

Parth (parthkohli):

Lol my smartscore is 99 I should say that to you I'm your senior ;P @ajprincess

OpenStudy (anonymous):

I'm still confused. I just don't get it.

Parth (parthkohli):

Practice with some problems. You'll get it ;p

OpenStudy (anonymous):

I guessed on that answer. I don't get the split part and I'm still alittle confused about the part where I find sum of b and product of c. or something.. I don't know..

OpenStudy (anonymous):

10x^2 + 2x − 8. 10 * -8 = -80. 10 - 8 = 2. OH WOW.

OpenStudy (anonymous):

2 = b.

OpenStudy (ajprincess):

u r right @careless850.

OpenStudy (anonymous):

But then what do I do?

Parth (parthkohli):

Don't confuse the variables in your mind. Never do that. How I learnt was just the multiplication of first and last term's coefficient and then expressing the middle term as a sum of two numbers that have their product equal the multiplication of first and last coefficients.

OpenStudy (anonymous):

Hm. So what do I do now? Because I have to turn them into the factorization.

Parth (parthkohli):

See the first two and last two terms. Factorize them and then group them. See what I have done in the answer very carefully.

OpenStudy (anonymous):

out of the options. --> 2(5x − 1)(x + 4) 2(5x − 4)(x + 1) 2(5x + 2)(x − 2) 2(5x − 2)(x + 2). which one could I eliminate right off the start?

Parth (parthkohli):

Lol if you have options, you can multiply all and check :P no need to learn factorization hehe

OpenStudy (anonymous):

I looked at the one you did earlier. Well, my next question don't have options.

OpenStudy (anonymous):

And I don't see why you put the x infront of the (3x-4)

OpenStudy (anonymous):

on the 2nd part.

Parth (parthkohli):

Then see my method again and again until you get it....or see khanacademy.

Parth (parthkohli):

http://www.youtube.com/watch?v=X7B_tH4O-_s

OpenStudy (ajprincess):

first u can take 2 out from the expression 10x^2 + 2x − 8 bcos 2 is common to all 10, 2 and 8. Is it clear?